OpenStudy (vespersaviour):

Which is the quadratic variation equation for the relationship? y varies directly with x^2 and y = 48 when x = 2. A. y = 12x^2 B. y = 4x^2 C. y = 4x D. y = x^2 + 25

OpenStudy (vespersaviour):

y=kx but for me y=kx^2 so plugging the numbers in 48=k(2)^2 48/4=12 ..so my guess is A? @MrNood can confirm?